OpenStudy (anonymous):

Find the minimum or maximum value of the function g(x) = -x2 + 3x - 2. Then state the domain and range of the function.

OpenStudy (raden):

just find the value of g(-b/2a) this case, given a=-1 and b=3 so, g(-3/2(-1)) = g(3/2) = ... ?

OpenStudy (raden):

btw, it is maximum because the coefficient of x^2 < 0

OpenStudy (anonymous):

all real numbers?

OpenStudy (raden):

yes!

OpenStudy (anonymous):

There are about 1,400,000 drops in 25 gallons of liquid. What percent of a gallon is a single drop?

OpenStudy (anonymous):

maximum value is 1/4 min. value does not exist. domain is all real no. range is less than or equal to 1/4.